Hi, I was wondering if someone could possibly help me. I am supposed to do these homework questions in my Algebra class, but I am having a hard time getting started. I don't know what steps to go through to obtain the answer. The back of the book shows the answers, but that doesn't help me much. I need to show I got to the answer. If anyone could possibly assist me, by showing me how you obtained the answer, I would be forever grateful! :o)
(Section 11.1, Pg. 749, # 107)
Directions: In the following exercise use the compound interest formula
A=P(1+r) ^t to find the annual interest rate, r.
In 2 years, an investment of $1280 grows to $1445.
the answer in the back of the book is: 6.25%, but I have no idea how they got this answer. Your help is forever appreciated! :o)

Answer by edjones(8007)

A=P(1+r)^t
1445=1280(1+r)^2
(1+r)^2=289/256
1+r=1.0625 square root of each side.
r=.0625=6.25%
.
Ed